Metallic sodium can be made by the electrolysis of molten NaCl.

(a) What mass of Na is formed by passing a current of 9.02 A through molten NaCl for 1.60 days? The unbalanced chemical reaction representing this electrolysis is shown below. NaCl Na + Cl2 g of Na is formed by this electrolysis.

(b) How many minutes are needed to plate out 6.00 g of Na from molten NaCl using 9.49 A current? minutes are needed.
Answer #1

2NaCl(aq) ---------> 2Na^+ (aq) + 2Cl^- (aq)

Na^+ (aq) + e^- -----------> Na(s)

z = 1

M = 23g/mole [ atomic mass of Na]

C   = 9.02A

t = 1.60 days = 1.6*24*60*60   = 138240seconds

W   = MCt/ZF

       = 23*9.02*138240/(1*96500)   = 297.2g

2.

Na^+ (aq) + e^- -----------> Na(s)

z = 1

M = 23g/mole [ atomic mass of Na]

C   = 9.49A

W = 6g

W    = MCt/ZF

t    = WZF/MC

      = 6*1*96500/(23*9.49)

      = 2652.67seconds

     = 2652.67/60

     = 44.2minutes >>>>answer